# Every Wrinkle Carries A Memory: An Integro-differential Bootstrap for Features in Cosmological Correlators

### v1

@@ 1  Introduction
The above diagram lives in exact de Sitter space and therefore is invariant under its full $SO(4,1)$ isometry group, which notably includes three de Sitter boosts and one dilation (a.k.a. scale invariance). However, realistic inflation turns out to be less symmetric and is thus crucially distinct from exact de Sitter. More specifically, the three de Sitter boosts are generically strongly broken during inflation since the rolling inflaton background picks a preferred rest frame, while the scale invariance is softly broken by the tilt of the inflaton potential. In fact, from established frameworks such as the EFT of inflation and its generalisations [38, 39, 40, 41, 42], we have learned that rich phenomenology, associated with sizeable non-Gaussianities, comes with a strong breaking of de Sitter isometries [43]. For instance, in cases where the inflaton kinetic energy driven by $\dot{\bar{\phi}}^{1/2}\sim 58H$ is pumped into the massive sector, the particle production rate can be exponentially enhanced. As a direct consequence of copious particle production, the cosmological collider signals, which encode the particle spectrum during inflation in the form of squeezed-limit non-analytic dependences on the momentum ratio [40, 44, 45, 46, 47], are dramatically amplified, alleviating their Boltzmann suppression from a de Sitter-invariant universe [48, 49, 50, 51, 52, 53, 54, 55, 11, 10, 56, 57, 58, 59, 60, 61]. Motivated by these observations, significant recent progress has been made in bootstrapping boostless correlators using non-symmetry-based tools. These include, for example, modified versions of the bootstrap differential equation (1.1) (e.g. in the presence of non-trivial sound speeds[14, 62, 15, 27]), cutting rules based on unitarity [63, 64, 65, 66, 67, 68] and causality [69, 70, 71, 72] , as well as analyticity & recursion relations [73, 74, 75, 76, 77].
@@
In this work, we shall leap away from the scale-invariance lamppost by proposing a bootstrap of general scale-breaking tree correlators. When de Sitter dilatation is broken, a local time evolution no longer translates into an ordinary local differential equation for exchange correlators. Instead, from correspondence between conformal time $\eta$ and comoving scale $k$ (i.e. $k=-1/\eta$, at horizon crossing [^3] Note that, despite the breaking of scale invariance in the bulk-bulk propagator, the conformal time-comoving scale correspondence is upheld by the scale-invariant bulk-boundary propagator $\propto\exp(ik\eta)$.), one should expect non-local relations to arise between widely separated configurations in the kinematic space. Such boundary non-localities resonate with the notion that correlators at the end of inflation encode an integrated history over the bulk, the details of which cannot be extracted from any one momentum configuration. Indeed, we will find that in cases where scale breaking solely originates from the massive sector (via explicit time dependence in either the free theory dynamics or interactions), the resulting four-point diagram satisfies an integro-differential equation of the schematic form [^4] More precisely, it is the Schwinger-Keldysh components $F_{\pm\pm}$ that satisfy integro differential equations like (1.2). See (1.10). {keyeqn}
@@
where we have omitted the functional dependence on $v$ for simplicity. The effect of scale breaking is carried by the memory kernel $K(x)$, whose detailed form relies on specific scale-breaking models, such as the heavy field’s mass or sound-speed time-dependence. But in all cases (under the Bunch-Davies initial condition assumption) we will show that $K(x)$ must identically vanish for $x<0$. As illustrated in Figure 1, this imposes a retardation in kinematic space for (1.2), where configurations with $u^{\prime}>u$ (less squeezed than the left hand side) are precluded from the right hand side. Also note that, given the explicit dependence of the kernel $K$ on the comoving momentum $s$, neither the integro-differential equation nor its solutions $F(u;s)$ are symmetric under the rescaling $\bm{k}_{i}\to\lambda\,\bm{k}_{i}$.
@@
where $\infty<\eta<0$ is the conformal time. We occasionally use the FLRW coordinates, ${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t){\rm d}\bm{x}^{2}$ with $a(t)=e^{Ht}$. To avoid clutter, we mostly set the Hubble rate to unity ($H=1$) unless otherwise stated. A heavy field in de Sitter with mass $m_{0}$ is characterised by the dimensionless mass-index $\mu=(m_{0}^{2}/H^{2}-9/4)^{1/2}$ which we assume to be positive. We introduce the following shorthand involving factors of the Euler gamma function:
@@
[…]
@@ 1.1 Bootstrap roadmap
The exchange of such a massive field sources a four-point function (see Figure 1) whose Schwinger-Keldysh integrals $F_{\sf ab}$ ($\sf a,b=\pm$) satisfy an integro-differential bootstrap equation [^6] Unlike its Schwinger-Keldysh components, the full correlator, $F=\sum_{\mathsf{a},\mathsf{b}=\pm 1}F_{\mathsf{a}\mathsf{b}}$, does not satisfy any individual equations. As a result, one may prefer to use the quartic wavefunction coefficient $\psi_{4}$ associated with the same exchange diagram, which turns out to be governed by the same equation as that of $F_{++}$. Nevertheless, we chose the Schwinger-Keldysh components because they can be more directly mapped to observables (e.g. the power spectrum and the bispectrum). See Section 4 for more details.
@@
is a scale-breaking memory kernel reflecting the time dependence of the mass. In the simplest case where the mass is a monochromatic oscillatory function of time, $\rho_{\omega^{\prime}}=g^{2}[\delta(\omega^{\prime}-\omega)+\delta(\omega^{\prime}+\omega)]$, the memory kernel reduces by essentially dropping the frequency integral. Such oscillatory masses are motivated by models of axion monodromy inflation, where the approximate shift symmetry of the inflaton field $\phi$ is broken down to the discrete subgroup $\phi\to\phi+2\pi nf$, where $f$ is the axion decay constant. This allows incorporating new operators into the kinetic term of the heavy field such as $g^{2}m_{0}^{2}\cos(\phi/f)\sigma^{2}$, which produces an oscillatory mass correction with $\omega=\dot{\bar{\phi}}/f$ upon setting the inflaton to its background $\phi=\bar{\phi}(t)$.
@@
    Bunch-Davies vacuum: The structure of the integro-differential equations (1.10) is also constrained by the Bunch-Davies initial condition [^7] Note that an asymptotic Bunch-Davies vacuum is well-defined under the mild assumption that mass grows slower than the exponential $e^{-Ht}$ in the far past. of the massive field. We will show that the domain of the momentum integral in (1.10) is accordingly restricted to $q>0$, which, loosely speaking, corresponds to four-point kinematics that are more squeezed than the left hand side configuration (see Figure 1 for an illustration).
@@ 2  Model setup
As was explained in the introduction, the backdrop of our study is a quasi-single field [44] scenario in which the approximate shift symmetry of the Goldstone boson $\pi$ typically assumed in the EFT of inflation is explicitly broken[38, 48, 90, 39]. For concreteness, we assume that the inflaton field is a canonical scalar interacting with an additional scalar field $\sigma$. The inflaton is assumed to be approximately massless while the scalar $\sigma$ is massive. In the unitary gauge (where $\pi=0$), the action of the system at leading order in derivatives and up to quadratic order in $\sigma$ and metric perturbations is given by [^8] Notice that tadpole terms in $\sigma$ are not allowed.
@@
$\qquad\qquad\qquad-\tilde{M}_{1}(t)\delta g^{00}\sigma-\tilde{M}_{2}(t)(\delta g^{00})^{2}\sigma-\tilde{M}_{3}(t)\delta g^{00}\sigma^{2}\Bigg]\,.$
@@
By combining the quadratic and cubic vertices in (2.6), one can construct single-exchange diagrams, with two or three massless external legs (see Figure 2). Let us consider vertices of the forms $\lambda_{1}(t)\dot{\pi}\sigma$, $\lambda_{2}(t)\pi\dot{\pi}\sigma$, $\lambda_{3}(t)\dot{\pi}^{2}\sigma$ or $\lambda_{4}(t)(\partial_{i}\pi)^{2}\sigma$, where the time-dependent coupling $\lambda_{i}(t)$ is specified in each case by $\tilde{M}_{1}(t)$ and $\tilde{c}(t)$. Diagrams with such vertices can be efficiently extracted from a set of seed four-point functions, defined as single-exchange diagrams with four external conformally coupled fields $\varphi$ and cubic vertices of the form $\lambda_{i}(t)\varphi^{2}\sigma$. Thanks to the simpler bulk-boundary propagator of the conformally coupled field, these exchange diagrams are analytically easier to handle than their massless counterparts. Most notably, as will be shown in the next section, these diagrams satisfy a set of integro-differential equations (IDEs), which are the generalisations of the bootstrap differential equations for de Sitter seed diagrams (i.e. with constant intermediate masses)[1]. Once these IDEs are solved, the resulting seed four-point functions can be mapped onto the power spectrum and the bispectrum of curvature perturbations. We will achieve this in Section 4.3 through a set of weight-shifting operators and by taking appropriate soft limits to reduce from the four-point to the three- and the two-point kinematics, along the lines of [106, 14, 15].
@@ 3.1 General vertices, sound speeds and masses
where $\rho_{\omega}$ ($\rho^{*}_{-\omega}$) is a dimensionless quantity. Inserting the above transformation into (3.6) and using the identities [^10] Perhaps an inspiring way to understand these identities is to take the analogy of integer power-shifting derivatives, which give e.g. $\partial_{k_{12}}e^{ik_{12}\eta}\sim\eta e^{ik_{12}\eta}$, and apply their generalisation to fractional derivatives, yielding $D^{\alpha}_{k_{12}}e^{ik_{12}\eta}\sim\int_{k_{12}}^{\infty}{\rm d}p\,(p-k_{12})^{-1-\alpha}e^{ip\eta}\sim\eta^{\alpha}e^{ik_{12}\eta}$, essentially giving rise to (3.8).
@@
we find an integral relation between $J_{\mathsf{a}\mathsf{b}}$ and $F_{\mathsf{a}\mathsf{b}}$. [^11] Note that the convergence of the time integrals requires the first and second lines to be substituted into $J_{+\pm}$ and $J_{-\pm}$, respectively. Plugging this relation back into (3.1) finally yields a closed set of integro-differential equations (IDEs) for $F_{\mathsf{a}\mathsf{b}}$: {keyeqn}
@@
A few remarks on our integro-differential equations are in order. First, we note that even though each Schwinger-Keldysh component $F_{\mathsf{a}\mathsf{b}}$ satisfies an IDE, the full correlator $F=\sum_{\mathsf{a}\mathsf{b}}F_{\mathsf{a}\mathsf{b}}$ does not due to the difference between the IDE kernels, i.e. $\Pi_{+\pm}\neq\Pi_{-\mp}$. Moreover, it may sound discouraging that our IDEs rely on the diagram’s Schwinger-Keldysh components, which are not individually well-defined observables on the boundary. To mitigate this issue, one may recast the problem in terms of the wavefunction coefficient $\psi_{4}$ associated with the same exchange graph. This quantity is in principle observable, with the advantage of satisfying an IDE identical to that obeyed by $F_{++}$. However, even after solving this IDE for $\psi_{4}$ one still needs the mode function (at the end of inflation) to extract practical observables such as the power spectrum and the bispectrum. [^12] The full correlator $F$ can indeed be expressed in terms of the wavefunction coefficients $\psi_{4}$ (for the exchange diagram) and $\psi_{3}$ (for its cubic contact sub-diagram), with the caveat that this relationship relies on the explicit form of the mode function $\sigma(s,\eta_{\rm end})$, the closed-form of which is unknown. This justifies our preference for working with the Schwinger-Keldysh components $F_{\mathsf{a}\mathsf{b}}$, allowing us to extract such observables without any direct reference to the explicit form of the mode function. See Section 4.3 for the derivations.
Second, we highlight again that the integro-differential equations (3.9) are intrinsically non-local in kinematic space only in a restricted sense: At fixed $k_{34}$ and $s$, the behaviour of $F_{\mathsf{a}\mathsf{b}}(k_{12})$ on the LHS is only tied to four-point configurations that are more squeezed on the RHS, i.e. those with $k^{\prime}_{12}=k_{12}+q>k_{12}$. This kinematic retardation is a direct avatar of the Bunch-Davies initial condition which is built into $F_{\mathsf{a}\mathsf{b}}$ via the $i\epsilon$ prescription. Indeed, in deriving the IDEs, the $i\epsilon$ prescription for $F_{+\pm}$($F_{-\pm}$) required decomposing the mass perturbation $\Delta m^{2}$ into negative (positive) frequency plane-waves $e^{+iq\eta}$ ($e^{-iq\eta}$), before inserting them into $J_{+\pm}$($J_{-\pm}$) in (3.6). Consequently, $F_{\mathsf{a}\mathsf{b}}(k_{12}+q,k_{34},s)$ always appears with $q>0$ in (3.9).

### v2

@@ 1  Introduction
The above diagram lives in exact de Sitter space and therefore is invariant under its full $SO(4,1)$ isometry group, which notably includes three de Sitter boosts and one dilation (a.k.a. scale invariance). However, realistic inflation turns out to be less symmetric and is thus crucially distinct from exact de Sitter. More specifically, the three de Sitter boosts are generically strongly broken during inflation since the rolling inflaton background picks a preferred rest frame, while the scale invariance is softly broken by the tilt of the inflaton potential. In fact, from established frameworks such as the EFT of inflation and its generalisations [38, 39, 40, 41, 42, 43], we have learned that rich phenomenology, associated with sizeable non-Gaussianities, comes with a strong breaking of de Sitter isometries [44]. For instance, in cases where the inflaton kinetic energy driven by $\dot{\bar{\phi}}^{1/2}\sim 58H$ is pumped into the massive sector, the particle production rate can be exponentially enhanced. As a direct consequence of copious particle production, the cosmological collider signals, which encode the particle spectrum during inflation in the form of squeezed-limit non-analytic dependences on the momentum ratio [40, 45, 46, 47, 48], are dramatically amplified, alleviating their Boltzmann suppression from a de Sitter-invariant universe [49, 50, 51, 52, 53, 54, 55, 56, 11, 10, 57, 58, 59, 60, 61, 62]. Motivated by these observations, significant recent progress has been made in bootstrapping boostless correlators using non-symmetry-based tools. These include, for example, modified versions of the bootstrap differential equation (1.1) (e.g. in the presence of non-trivial sound speeds[14, 63, 15, 27]), cutting rules based on unitarity [64, 65, 66, 67, 68, 69] and causality [70, 71, 72, 73] , as well as analyticity & recursion relations [74, 75, 76, 77, 78].
@@
In this work, we shall leap away from the scale-invariance lamppost by proposing a bootstrap of general scale-breaking tree correlators. When de Sitter dilatation is broken, a local time evolution no longer translates into an ordinary local differential equation for exchange correlators. Instead, from the correspondence between conformal time $\eta$ and comoving scale $k$ (i.e. $k=-1/\eta$, at horizon crossing [^3] Note that, despite the breaking of scale invariance in the bulk-bulk propagator, the conformal time-comoving scale correspondence is upheld by the scale-invariant bulk-boundary propagator $\propto\exp(ik\eta)$.), one should expect non-local relations to arise between widely separated configurations in kinematic space. Such boundary non-localities resonate with the notion that correlators at the end of inflation encode an integrated history over the bulk, whose details cannot be extracted from any one momentum configuration. Indeed, we will find that in cases where scale breaking solely originates from the massive sector (via explicit time dependence in either the free theory dynamics or interactions), the resulting four-point diagram satisfies an integro-differential equation of the schematic form [^4] More precisely, it is the Schwinger-Keldysh components $F_{\pm\pm}$ that satisfy integro differential equations like (1.2). See (1.10). {keyeqn}
@@
where we have omitted the functional dependence on $v$ for simplicity. The effect of scale breaking is carried by the memory kernel $K(x)$, whose detailed form relies on specific scale-breaking models, such as the heavy field’s mass or sound-speed time-dependence. But in all cases (under the Bunch-Davies initial condition assumption) we will show that $K(x)$ must identically vanish for $x<0$. As illustrated in Figure 1, this imposes a retardation in kinematic space for (1.2), where configurations with $u>u^{\prime}$ (less squeezed than the left hand side) are precluded from the right hand side. Also note that, given the explicit dependence of the kernel $K$ on the comoving momentum $s$, neither the integro-differential equation nor its solutions $F(u;s)$ are symmetric under the rescaling $\bm{k}_{i}\to\lambda\,\bm{k}_{i}$.
@@
where $\infty<\eta<0$ is the conformal time. We occasionally use the FLRW coordinates, ${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t){\rm d}\bm{x}^{2}$ with $a(t)=e^{Ht}$. To avoid clutter, we mostly set the Hubble rate to unity ($H=1$) unless otherwise stated. In the main body of the paper, we will consider a heavy bulk field with mass $m_{0}$ that is also characterised by the dimensionless mass-index $\mu=(m_{0}^{2}/H^{2}-9/4)^{1/2}>0$. The case with a light bulk field is discussed separately in Appendix C. We introduce the following shorthand involving factors of the Euler gamma function:
@@
as convenient bootstrap variables. $\sigma$ will denote a heavy field whose mass oscillates in time as $m^{2}(t)=m_{0}^{2}+g^{2}m_{0}^{2}\cos[\omega(t-t_{0})]$, where $0\leq g<1$. We use the dimensionless parameter $x_{0}\equiv-s\eta_{0}$, in which $\eta_{0}=-\exp(-Ht_{0})$, as a proxy for the oscillation phase (note that $\eta_{0}$ is not to be confused with the boundary time $\eta_{\rm end}$ at the end of inflation). The power spectrum of the curvature perturbation is defined as
@@ 1.1 Bootstrap roadmap
The exchange of such a massive field sources a four-point function (see Figure 1) whose Schwinger-Keldysh integrals $F_{\sf ab}$ ($\sf a,b=\pm$) satisfy an integro-differential bootstrap equation
@@
is a scale-breaking memory kernel reflecting the time dependence of the mass. As an aside, note that the full correlator, $F=\sum_{\mathsf{a},\mathsf{b}=\pm 1}F_{\mathsf{a}\mathsf{b}}$, does not satisfy on its own any integro-differential equation like Eq. (1.10). The reason is the use of different memory kernels for individual Schwinger-Keldysh components. More details on this point will be provided in Section 3.
In the simplest case where the mass is a monochromatic oscillatory function of time, $\rho_{\omega^{\prime}}=g^{2}[\delta(\omega^{\prime}-\omega)+\delta(\omega^{\prime}+\omega)]$, the memory kernel reduces by essentially dropping the frequency integral. Such oscillatory masses are motivated by models of axion monodromy inflation, where the approximate shift symmetry of the inflaton field $\phi$ is broken down to the discrete subgroup $\phi\to\phi+2\pi nf$, where $f$ is the axion decay constant. This allows incorporating new operators into the kinetic term of the heavy field such as $g^{2}m_{0}^{2}\cos(\phi/f)\sigma^{2}$, which produces an oscillatory mass correction with $\omega=\dot{\bar{\phi}}/f$ upon setting the inflaton to its background $\phi=\bar{\phi}(t)$.
@@
    Bunch-Davies vacuum: The structure of the integro-differential equations (1.10) is also constrained by the Bunch-Davies initial condition of the massive field. Such asymptotic vacuum is well-defined under the mild assumption that mass grows slower than the exponential $e^{-Ht}$ in the far past. We will show that, in (1.10), the domain of the momentum integral, due to this vacuum choice, is restricted to $q>0$, which, loosely speaking, corresponds to four-point kinematics that are more squeezed than the left hand side configuration (see Figure 1 for an illustration).
@@ 2  Model setup
As was explained in the introduction, the backdrop of our study is a quasi-single field [45] scenario in which the approximate shift symmetry of the Goldstone boson $\pi$ typically assumed in the EFT of inflation is explicitly broken[38, 49, 91, 39]. For concreteness, we assume that the inflaton field is a canonical scalar interacting with an additional scalar field $\sigma$. The inflaton is assumed to be approximately massless while the scalar $\sigma$ is massive. In the unitary gauge (where $\pi=0$), the action of the system at leading order in derivatives and up to quadratic order in $\sigma$ and metric perturbations is given by
@@
$\qquad\qquad\qquad-\tilde{M}_{1}(t)\delta g^{00}\sigma-\tilde{M}_{2}(t)(\delta g^{00})^{2}\sigma-\tilde{M}_{3}(t)\delta g^{00}\sigma^{2}\Bigg]\,$,
where tadpole terms are excluded.
@@
By combining the quadratic and cubic vertices in (2.6), one can construct single-exchange diagrams, with two or three massless external legs (see Figure 2). Let us consider vertices of the forms $\lambda_{1}(t)\dot{\pi}\sigma$, $\lambda_{2}(t)\pi\dot{\pi}\sigma$, $\lambda_{3}(t)\dot{\pi}^{2}\sigma$ or $\lambda_{4}(t)(\partial_{i}\pi)^{2}\sigma$, where the time-dependent coupling $\lambda_{i}(t)$ is specified in each case by $\tilde{M}_{1}(t)$ and $\tilde{c}(t)$. Diagrams with such vertices can be efficiently extracted from a set of seed four-point functions, defined as single-exchange diagrams with four external conformally coupled fields $\varphi$ and cubic vertices of the form $\lambda_{i}(t)\varphi^{2}\sigma$. Thanks to the simpler bulk-boundary propagator of the conformally coupled field, these exchange diagrams are easier to handle analytically than their massless counterparts. Most notably, as will be shown in the next section, these diagrams satisfy a set of integro-differential equations (IDEs), which are the generalisations of the bootstrap differential equations for de Sitter seed diagrams (with constant intermediate masses)[1]. Once these IDEs are solved, the resulting seed four-point functions can be mapped onto the power spectrum and the bispectrum of curvature perturbations. We will achieve this in Section 4.3 through a set of weight-shifting operators and by taking appropriate soft limits to reduce from the four-point to the three- and the two-point kinematics, along the lines of [107, 14, 15].
@@ 3.1 General vertices, sound speeds and masses
where $\rho_{\omega}$ ($\rho^{*}_{-\omega}$) is a dimensionless quantity. Inserting the above transformation into (3.6) and using the identities
@@
we find an integral relation between $J_{\mathsf{a}\mathsf{b}}$ and $F_{\mathsf{a}\mathsf{b}}$. Note that the convergence of the time integrals requires the first and second lines to be substituted into $J_{+\pm}$ and $J_{-\pm}$, respectively. An inspiring way to understand these identities is to take the analogy of integer power-shifting derivatives, which give e.g. $\partial_{k_{12}}e^{ik_{12}\eta}\sim\eta e^{ik_{12}\eta}$, and apply their generalisation to fractional derivatives, yielding $D^{\alpha}_{k_{12}}e^{ik_{12}\eta}\sim\int_{k_{12}}^{\infty}{\rm d}p\,(p-k_{12})^{-1-\alpha}e^{ip\eta}\sim\eta^{\alpha}e^{ik_{12}\eta}$, essentially giving rise to (3.8).
Plugging this relation back into (3.1) finally yields a closed set of integro-differential equations (IDEs) for $F_{\mathsf{a}\mathsf{b}}$: {keyeqn}
@@
A few remarks on our integro-differential equations are in order. First, we note that unlike its Schwinger-Keldysh components $F_{\mathsf{a}\mathsf{b}}$, the full correlator $F=\sum_{\mathsf{a}\mathsf{b}}F_{\mathsf{a}\mathsf{b}}$ does not satisfy any IDE. This is due to the difference between the IDE kernels, i.e. $\Pi_{+\pm}\neq\Pi_{-\mp}$. In fact, it may seem discouraging that our IDEs are based on individual Schwinger-Keldysh components of the graph, which are not well-defined observables in isolation. As an alternative, one may choose to work with the wavefunction coefficient $\psi_{4}$ associated with the same exchange graph, with the advantage of being an observable that satisfies an IDE identical to that of $F_{++}$. However, even after solving the IDE for $\psi_{4}$, one would still require the mode function at $\eta_{\text{emd}}$ in order to compute relevant boundary observables at the end of inflation, such as the power spectrum and the bispectrum. This justifies our preference for working with the Schwinger-Keldysh components $F_{\mathsf{a}\mathsf{b}}$, allowing us to directly extract the full correlator without any explicit reference to the mode function $\sigma(s,\eta_{\rm end})$, the closed-form of which is unknown. See Section 4.3 for the derivations.
Second, we highlight again that the integro-differential equations (3.9) are intrinsically non-local in kinematic space only in a restricted sense: At fixed $k_{34}$ and $s$, the behaviour of $F_{\mathsf{a}\mathsf{b}}(k_{12})$ on the LHS is only tied to four-point configurations that are more squeezed on the RHS, i.e. those with $k^{\prime}_{12}=k_{12}+q>k_{12}$. This kinematic retardation is a direct avatar of the Bunch-Davies initial condition which is built into $F_{\mathsf{a}\mathsf{b}}$ via the $i\epsilon$ prescription. Indeed, in deriving the IDEs, the $i\epsilon$ prescription for $F_{+\pm}$($F_{-\pm}$) required us to decompose the mass perturbation $\Delta m^{2}$ into negative (positive) frequency plane-waves $e^{+iq\eta}$ ($e^{-iq\eta}$), before inserting them into $J_{+\pm}$($J_{-\pm}$) in (3.6). Consequently, $F_{\mathsf{a}\mathsf{b}}(k_{12}+q,k_{34},s)$ always appears with $q>0$ in (3.9).